# Contraction Properties of the Global Workspace Primitive

### v1

@@ 1.1 Global Workspace Theory
Besides studying modular structures that result from different training paradigms, we can take greater inspiration from known neuroanatomical properties to impose a specific multi-area topology on an "RNN of RNNs", thereby investigating how a given structural constraint impacts various facets of model performance. One such structure of special salience is derived from the global workspace (GW) theories of cognition [6, 7, 8]. The GW combination is composed of a central subnetwork through which all other subnetworks interact (Figure 1). The core idea of the global workspace is that it serves as a shared latent space for communication across the network. It is a prefrontal cortex-like structure that in a human cognitive interpretation represents conscious experience and sensory imagination.
@@ 1.2 Stability in Theory and Practice
One of the largest gaps in stability theory for RNNs at present reflects the broader paucity of theoretical neuroscience work on multi-area recurrent networks. To our knowledge, the only work to date on stability of "RNNs of RNNs" is that of Kozachkov et al. [1], which utilizes contractive stability. Loosely speaking, a dynamical system is said to be globally contracting if given any two system trajectories, they will converge to each other exponentially in a valid distance metric, regardless of initial conditions [32]. A major benefit of contractive stability is that there are a number of ways in which contracting systems can be combined whilst preserving stability of the overall system. Although the importance of building modular systems is clear, in general there is no guarantee that a combination of stable systems will itself be stable. Because contraction analysis tools allow complicated contracting systems to be built up recursively from simpler elements, this form of stability is well suited for neural and other biological systems [29]. Moreover, contracting combinations can be made between systems with very different dynamics, as long as those dynamics are each contracting – thereby facilitating a great deal of flexibility in constructing stable modular systems [33, 34].
@@ 2 The "Stable RNN of RNNs" Model Class
Throughout this paper we use a multi-area recurrent neural network model, i.e. an "RNN of RNNs", as introduced with stability conditions in [1]. To summarize, each individual RNN used as a subnetwork is described by a well-established continuous-time nonlinear RNN model equation [38], and then the units within those different individual RNNs are connected across the subnetwork "areas" via a different component of the overall model. To construct stable networks of networks, we perform these inter-area combinations in a way that keeps the overall system dynamics stable so long as the subnetwork dynamics are independently stable, via contraction analysis [34]. Unfortunately, there is minimal prior work on stability of "RNNs of RNNs", so we focus our mathematical background on explaining the results of Kozachkov et al. [1]. For additional discussion of stability conditions for individual RNNs, see the introduction in [1].
@@ 2.1 Model Definition
${\mathbf{x}}=\mathbf{f}(\mathbf{x},t)=-\mathbf{x}+\mathbf{W}\phi(\mathbf{x})+\mathbf{u}(t)+\mathbf{L}\mathbf{x}\ \in\mathbb{R}^{N}$ (overall ”RNN of RNNs” dynamics)
@@ 3 Mathematical Results
In sum, because of the wide ranging opportunities for expansion on the theoretical results of Kozachkov et al. [1], our mathematical results largely involve relaxation of certain stability conditions originally reported there. By considering constraints on inter-area topologies and inter-area weight magnitudes instead, we avoid the restrictiveness of linear pure negative feedback. Specifically, we prove a more permissive multi-area network contraction condition in the special case of a global workspace modular topology (subsection 3.1). We also close with mathematical remarks linked to our current experimental aims on the discovery of effective sparse multi-area connection topologies (subsection 3.2).
@@ 3.1 Global Workspace Combinations
${\mathbf{x}}=\mathbf{f}(\mathbf{x},t)=-\mathbf{x}+\mathbf{W}\phi(\mathbf{x})+\mathbf{u}(t)+\mathbf{L}\psi(\mathbf{x})\ \in\mathbb{R}^{N}$ (global workspace ”RNN of RNNs” dynamics)
@@
For the architecture constrained to have GW topology, the inter-area connectivity matrix $\mathbf{L}$ takes on a special structure, which can be exploited in proofs of stability and learning convergence. We adopt the convention that the ‘central’ subnetwork (Figure 1) is denoted by index 0. Then the $i,j$ block of matrix $\mathbf{L}$ ($L_{ij}$, corresponding to the weights leaving subnetwork $j$ and arriving at subnetwork $i$) are zero unless $i=0$ or $j=0$. Note that the diagonal block $i=0,\ j=0$ is zero by definition, because $\mathbf{L}$ is restricted to only connections between distinct subnetworks. Given this global workspace structure for $\mathbf{L}$, along with subnetworks known to be individually contracting in some identifiable metric, we now set out to find stability conditions for the resulting system with dynamics that evolve according to (global workspace ”RNN of RNNs” dynamics). We can also think of this combination property as utilizing repeated nonlinear couplings, to ultimately create a flower-like structure.
@@
We are interested in using the spectral norm of each such block. Because the spectral norm of a matrix and its transpose are equivalent, this reduces our consideration to the max eigenvalues for $p-1$ total unique matrices, defined as the symmetric part of each above off-diagonal block. We thus have for $i=1$ to $p$:
@@
$||L_{i0}||\leq\ell\hskip 28.45274pt\text{i.e.}\hskip 28.45274pt\frac{L_{i0}+L_{i0}^{T}}{2}\preceq\ell$
@@
$||G_{i}||\leq g_{\psi}\ell$
Using the GW block structure imposed on $(\mathbf{L}\mathbf{D}_{\psi})_{sym}$ and the bounds on the spectral norms of all relevant blocks $||G_{i}||$, we can now ensure that the symmetric part of the overall system Jacobian is negative semi-definite via the Schur Complement Theorem [42]. This yields the following sufficient condition for contraction of the overall system in terms of $\ell$, the small-gain criterion for the inter-area weights as desired:
@@
Like the negative feedback described by Kozachkov et al. [1], the GW combination property can be used as a primitive for recursive construction of "RNNs of RNNs of RNNs (…)". This enables a number of interesting structures to be built at scale, including e.g. a GW combination of multiple all-to-all negative feedback multi-area networks.
@@
Note that in the special case where the inter-area connections are linear (i.e. $\psi(x)=x$), the constraint on $||L_{ij}||$ need not scale with the number of subnetworks $p$. This is because for linear connections $\mathbf{D_{\psi}}=\mathbf{I}$, so skew-symmetry (with respect to a metric) leads to perfect cancellation and thus does not depend on $p$. One could also consider a model with two different types of inter-areal connections between subnetworks, one linear and one nonlinear – then only the nonlinear connectivity matrix should scale with the number of subnetworks $p$.
@@ 3.2 Theoretical Context for Experimental Aims
Our pilot empirical work to be presented in the subsequent section primarily utilizes the model (overall ”RNN of RNNs” dynamics) with nonlinear subnetwork stability enforced by the weight magnitudes condition in Theorem 1 of [1] and overall model stability enforced via the condition (linear negative feedback parameterization). However as the discussion by Kozachkov et al. [1] primarily assumes negative feedback will be perfect and all-to-all, there are a few important points to make before exploring variations to the inter-area structure. Particularly in the context of removal of weights or units from the network, as these are topics of high empirical priority, but in general stability preservation cannot be guaranteed with such actions.
@@ 4 Experimental Results
Broadly, we have a few different categories of major empirical questions about the stable "RNNs of RNNs" model class stemming from the pilot work by Kozachkov et al. [1]:
@@ 4.1 Methods Summary
Like in [1], we evaluated different provably stable "network of networks" architectures using sequence learning benchmark tasks. Kozachkov et al. [1] focused on three sequential image classification tasks often used to evaluate RNN performance: sequential MNIST, permuted sequential MNIST, and sequential CIFAR10. These tasks measure information processing ability over long sequences [46], and are particularly common to use in benchmarking provably stable RNNs, as stability in theory has a tradeoff with the ability to integrate information over long timescales – though in practice stable RNNs have largely performed well with sequence processing [5]. For our experiments, we used the ’Sparse Combo Net’ code provided by [1], and performed evaluations using the hardest sequential CIFAR10 (seqCIFAR10) task. Therefore the network’s dynamics were defined according to a discrete-time version of equation (overall ”RNN of RNNs” dynamics), with ϕ=italic-$\phi=$ ReLU and $\mathbf{L}$ parameterized according to (linear negative feedback parameterization). We also retain the ’$p\times n$ network’ notation of [1] in reporting our results; such a network consists of $p$ distinct subnetwork RNNs, with each such subnetwork RNN containing $n$ units.
@@ 4.3 Task Performance Varies with Negative Feedback Structure
Name
Trainable Params Count
seqCIFAR10
@@
10x32 All-to-All Negative Feedback
50,570
62.78%
0.9198
16x32 Global Workspace
22,538
57.92%
1.0921
16x32 All-to-All Negative Feedback
130,058
66.85%
0.765
24x32 Random 5% of Negative Feedback Pairings
24,074
34.23%
7.9304
24x32 Global Workspace
34,313
62.59%
0.8161
24x32 Random 50% of Negative Feedback Pairings
142,858
68.91%
0.6062
24x32 All-to-All Negative Feedback
293,386
53.62%
1.3028
32x32 Global Workspace
46,090
62.5%
0.7904
32x32 Random 33.3% of Negative Feedback Pairings
183,306
68.64%
0.541
@@ 4.5 Variations on Subnetwork Components
Name
Trainable Params Count
seqCIFAR10
@@
16x32 Global Workspace Included
22,538
57.92%
1.0921
24x32 Global Workspace Included
34,313
62.59%
0.8161
24x32 Global Workspace Repetition
34,313
61.52%
0.8644
24x32 Global Workspace Selection
34,313
62.32%
0.841
Size 64 Global Workspace With 24x32 Subnets
60,810
63.72%
0.6806
32x32 Global Workspace
46,090
62.5%
0.7904
Size 128 Global Workspace With 32x32 Subnets
147,210
65.57%
0.5532
Table 2: Performance of Global Workspace Sparse Combo Nets with different central module settings. The same stats as in Table 1 are presented here, now for an expanded set of global workspace experiments. This includes the initial tests where an otherwise normal node in the Sparse Combo Net was made to be the GW, alongside any repetitions run (Figure S2). Additionally, two new trials (highlighted rows) are reported – a size 64 central module connected to a $24\times 32$ set of contributing subnetworks initialized as before, and a size 128 central module connected to a $32\times 32$ set of contributing subnetworks initialized as before. In both cases the separate global workspace subnetwork was initialized using a $1\%$ sparsity level and weights between $-5$ and $5$. The larger GW subnetworks both resulted in test accuracy improvements, though they had notably larger size in terms of trainable parameters, and were perhaps slightly overfitted when comparing their training losses to those found in Table 1.
Note that while we used the notation $p\times n$ to refer to different size Sparse Combo Nets through most of this text for convenience, there is of course no requirement for network sizes to be the same, and subnetworks of a variety of sizes can be just as readily described within the existing model definition. Overall, the Size 128 Global Workspace with $32\times 32$ Subnets matched the performance of the best network from [1], and reached this accuracy within the first 200 epochs of training. Both networks had a similar total number of trainable parameters (Tables 2 and S1), though the GW network had many more units and thus has the greater potential for expansion to other applications. The size 128 GW network was also quite resistant to post-training subnetwork ablation (Figure S4), unlike the $16\times 32$ Sparse Combo Net with all-to-all negative feedback introduced by Kozachkov et al. [1] was (Figure 2).
@@ 5.1 Future Directions
Note that task input at each timestep was first passed through a (trainable) linear input layer, represented by a weight matrix of size 3 (for RGB CIFAR pixels) by $N$, where $N=p*n$ is the total number of units in the $p\times n$ Combo RNN. On the flip side, the outputs of the Combo RNN neurons at the final timestep in the sequence were used for prediction via passing them through a (trainable) linear output layer. The output layer was represented by a matrix of size $N$ by 10, with 10 being the number of different classes contained in the CIFAR10 dataset. This allowed for computation of cross entropy loss during training based on the true class of each input. At test time, the output index with the maximum value was selected as the predicted label.
@@
Name
Stable RNN?
Params
sMNIST
@@
LSTM [49]
68K
97.3%
92.7%
59.7%
Transformer [50]
0.5M
98.9%
97.9%
62.2%
Antisymmetric [49]
36K
98%
95.8%
58.7%
Sparse Combo Net
130K
96.85% (4)
@@
As described in the main text, all experiments were run using the open sourced Sparse Combo Net code from [1] on Google Colab Pro+. Methodology for nonlinear subnetwork initialization and corresponding metric calculation was thus identical to the best performing trials of [1], as was the base model definition and general training loop. These settings were the same for all trials described in the main text, unless noted otherwise there. Across the whole paper (and all of [1]), Sparse Combo Net was run with $\tau=1$ and $dt=0.03$, such that after each timestep (i.e. after each pixel was presented), the current state was updated to be the previous state $+0.03*f(x)$, where $f(x)=-x+\mathbf{W}*\phi(x)+\mathbf{L}*x+u(t)$ with $x$ corresponding to the previous state of the neurons and $u(t)$ corresponding to the values returned by the linear input layer when given the pixel that was just input to the larger network. $\mathbf{W}$ and $\mathbf{L}$ are of course the intra- and inter- subnetwork weights of (overall ”RNN of RNNs” dynamics) respectively, so they are static throughout this process of evaluating an individual image sequence. Recall that before each new sequence, network starting state was reset.
@@
Like discussed within Table 2, we did use different initialization settings for the global workspace subnetwork when experimenting with differing sizes for this central module. The main difficulty in comparing subnetworks of different sizes for the Sparse Combo Net in general is that sparsity levels cannot be fairly compared – for a larger nonlinear RNN subnetwork it may be impossible to meet the Theorem 1 condition of [1] without either increasing sparsity or greatly decreasing allowed entry magnitude, but for a smaller nonlinear RNN subnetwork a highly sparse structure might mean that the majority of units have 0 non-zero weights and thus cannot meaningfully contribute. Indeed, there was a point at which increasing subnetwork sparsity decreased the performance of Sparse Combo Nets [1]. As such, we used different sparsity level for initialization of those larger global workspace subnetworks in our GW size experiments. Specifically, we tested a GW with 64 units attached to 24 typically-defined 32 unit subnetworks, and a GW with 128 units attached to 32 typically-defined 32 unit subnetworks. In both cases, the GW subnetwork was initialized with a $1\%$ sparsity level and entry magnitude between $-5$ and $5$. The networks were otherwise trained as before, and their performance on seqCIFAR10 (Table 2) was then compared to similarly sized Global Workspace Sparse Combo Nets from the previous trials, which would have GW subnetwork of size 32 with $3.3\%$ sparsity level and entry magnitude between $-6$ and $6$ (as was used for the Sparse Combo Nets more broadly in [1]).
One additional technical consideration that may warrant further testing in different network topologies is the initialization scheme for inter-area connections. Like [1], we sample each entry from a normal distribution with mean of 0 and standard deviation $\frac{1}{\sqrt{n}}$ where $n$ is the number of neurons in each subnetwork of a $p\times n$RNN of RNNs". In the case where different subnetworks have different sizes, this instead uses the mean number of neurons per subnetwork. With a smaller number of between-subnetwork connections (as in the global workspace topology), one might consider allowing nonzero entries to be initialized to higher entry magnitudes. While in the present paper we have made the most direct comparison possible, this adjustment could allow advantages of the more sparsely connected negative feedback architectures to be even better utilized [^1] For a more extended set of relevant future experimental directions than provided in this paper, see Ennis [43]..
@@
Name
Trainable Params Count
seqCIFAR10
@@
37,258
61.07%
0.93
Size 64 Global Workspace With 24x32 Subnets
60,810
63.72%
0.6806
2x32 Global Workspace With 24x32 Subnets
61,834
63.31%
0.7529
Size 128 Global Workspace With 32x32 Subnets
147,210
65.57%
0.5532
4x32 Global Workspace With 32x32 Subnets
153,354
64.75%
0.6546

### v2

@@ 1.1 Global Workspace Theory
Besides studying modular structures that result from different training paradigms, we can take greater inspiration from known neuroanatomical properties to impose a specific multi-area topology on an "RNN of RNNs", thereby investigating how a given structural constraint impacts various facets of model performance. One such structure of special salience is derived from the global workspace (GW) theories of cognition [6, 7, 8, 20, 21]. The GW combination is composed of a central subnetwork through which all other subnetworks interact (Figure 1). The core idea of the global workspace is that it serves as a shared latent space for communication across the network. It is a prefrontal cortex-like structure that in a human cognitive interpretation represents conscious experience and sensory imagination.
@@ 1.2 Stability in Theory and Practice
One of the largest gaps in stability theory for RNNs at present reflects the broader paucity of theoretical neuroscience work on multi-area recurrent networks [1]. Loosely speaking, a dynamical system is said to be globally contracting if given any two system trajectories, they will converge to each other exponentially in a valid distance metric, regardless of initial conditions [34]. A major benefit of contractive stability is that there are a number of ways in which contracting systems can be combined whilst preserving stability of the overall system. Although the importance of building modular systems is clear, in general there is no guarantee that a combination of stable systems will itself be stable. Because contraction analysis tools allow complicated contracting systems to be built up recursively from simpler elements, this form of stability is well suited for neural and other biological systems [31]. Moreover, contracting combinations can be made between systems with very different dynamics, as long as those dynamics are each contracting – thereby facilitating a great deal of flexibility in constructing stable modular systems [35, 36].
@@ 2 The "Stable RNN of RNNs" Model Class
Throughout this paper we use a multi-area recurrent neural network model, i.e. an "RNN of RNNs", as introduced with stability conditions in [1]. To summarize, each individual RNN used as a subnetwork is described by a well-established continuous-time nonlinear RNN model equation [40], and then the units within those different individual RNNs are connected across the subnetwork "areas" via a different component of the overall model. To construct stable networks of networks, we perform these inter-area combinations in a way that keeps the overall system dynamics stable so long as the subnetwork dynamics are independently stable, via contraction analysis [36]. Unfortunately, there is minimal prior work on stability of "RNNs of RNNs", so we focus our mathematical background on explaining previous results [1]. For additional discussion of stability conditions for individual RNNs, see the introduction in [1].
@@ 2.1 Model Definition
[…]
@@ 3 Mathematical Results
In sum, because of the wide ranging opportunities for expansion on previous theoretical results [1], our mathematical results largely involve relaxation of certain stability conditions originally reported there. By considering constraints on inter-area topologies and inter-area weight magnitudes instead, we avoid the restrictiveness of linear pure negative feedback. Specifically, we prove a more permissive multi-area network contraction condition in the special case of a global workspace modular topology (subsection 3.1). We also close with mathematical remarks linked to our current experimental aims on the discovery of effective sparse multi-area connection topologies (subsection 3.2).
@@ 3.1 Global Workspace Combinations
[…]
@@
For the architecture constrained to have GW topology, the inter-area connectivity matrix $\mathbf{L}$ takes on a special structure, which can be exploited in proofs of stability and learning convergence. We adopt the convention that the ‘central’ subnetwork (Figure 1) is denoted by index 0. Then the $i,j$ block of matrix $\mathbf{L}$ ($L_{ij}$, corresponding to the weights leaving subnetwork $j$ and arriving at subnetwork $i$) are zero unless $i=0$ or $j=0$. Note that the diagonal block i=0,j=0i=0,\ j=0 is zero by definition, because $\mathbf{L}$ is restricted to only connections between distinct subnetworks. Given this global workspace structure for $\mathbf{L}$, along with subnetworks known to be individually contracting in some identifiable metric, we now set out to find stability conditions for the resulting system with dynamics that evolve according to (global workspace ”RNN of RNNs” dynamics). We can also think of this combination property as utilizing repeated nonlinear couplings, to ultimately create a flower-like structure.
@@
We are interested in using the spectral norm of each such block. Because the spectral norm of a matrix and its transpose are equivalent, this reduces our consideration to the max eigenvalues for p1p-1 total unique matrices, defined as the symmetric part of each above off-diagonal block. We thus have for $i=1$ to $p$:
@@
$L_{i0}||\leq\ell\hskip 28.45274pt\text{i.e.}\hskip 28.45274pt\frac{L_{i0}+L_{i0}^{T}}{2}\preceq\ell$
@@
$G_{i}||\leq g_{\psi}\ell$
Using the GW block structure imposed on $(\mathbf{L}\mathbf{D}_{\psi})_{sym}$ and the bounds on the spectral norms of all relevant blocks $G_{i}||$, we can now ensure that the symmetric part of the overall system Jacobian is negative semi-definite via the Schur Complement Theorem [44]. This yields the following sufficient condition for contraction of the overall system in terms of $\ell$, the small-gain criterion for the inter-area weights as desired:
@@
The GW combination property can be used in a similar way to negative feedback as a primitive for recursive construction of "RNNs of RNNs of RNNs (…)". This enables a number of interesting structures to be built at scale, including e.g. a GW combination of multiple all-to-all negative feedback multi-area networks.
@@
Note that in the special case where the inter-area connections are linear (i.e. $\psi(x)=x$), the constraint on $L_{ij}||$ need not scale with the number of subnetworks $p$. This is because for linear connections $\mathbf{D_{\psi}}=\mathbf{I}$, so skew-symmetry (with respect to a metric) leads to perfect cancellation and thus does not depend on $p$. One could also consider a model with two different types of inter-areal connections between subnetworks, one linear and one nonlinear – then only the nonlinear connectivity matrix should scale with the number of subnetworks $p$.
@@ 3.2 Theoretical Context for Experimental Aims
Our pilot empirical work to be presented in the subsequent section primarily utilizes the model (overall ”RNN of RNNs” dynamics) with nonlinear subnetwork stability enforced by the weight magnitudes condition in Theorem 1 of [1] and overall model stability enforced via the condition (linear negative feedback parameterization). However previous work primarily assumes negative feedback will be perfect and all-to-all, there are a few important points to make before exploring variations to the inter-area structure [1]. Particularly in the context of removal of weights or units from the network, as these are topics of high empirical priority, but in general stability preservation cannot be guaranteed with such actions.
@@ 4 Experimental Results
Broadly, we have a few different categories of major empirical questions about the stable "RNNs of RNNs" model class stemming from previous work [1]:
@@ 4.1 Methods Summary
Like in [1], we evaluated different provably stable "network of networks" architectures using sequence learning benchmark tasks. Previous work [1] focused on three sequential image classification tasks often used to evaluate RNN performance: sequential MNIST, permuted sequential MNIST, and sequential CIFAR10. These tasks measure information processing ability over long sequences [48], and are particularly common to use in benchmarking provably stable RNNs, as stability in theory has a tradeoff with the ability to integrate information over long timescales – though in practice stable RNNs have largely performed well with sequence processing [5]. For our experiments, we used the ’Sparse Combo Net’ code provided by [1], and performed evaluations using the hardest sequential CIFAR10 (seqCIFAR10) task. Therefore the network’s dynamics were defined according to a discrete-time version of equation (overall ”RNN of RNNs” dynamics), with $\phi=$ ReLU and $\mathbf{L}$ parameterized according to (linear negative feedback parameterization). We also retain the $p\times n$ network’ notation of [1] in reporting our results; such a network consists of $p$ distinct subnetwork RNNs, with each such subnetwork RNN containing $n$ units.
@@ 4.3 Task Performance Varies with Negative Feedback Structure
Name Trainable Params Count seqCIFAR10
@@
10x32 All-to-All Negative Feedback 50,570 62.78% 0.9198
16x32 Global Workspace 22,538 57.92% 1.0921
16x32 All-to-All Negative Feedback 130,058 66.85% 0.765
24x32 Random 5% of Negative Feedback Pairings 24,074 34.23% 7.9304
24x32 Global Workspace 34,313 62.59% 0.8161
24x32 Random 50% of Negative Feedback Pairings 142,858 68.91% 0.6062
24x32 All-to-All Negative Feedback 293,386 53.62% 1.3028
32x32 Global Workspace 46,090 62.5% 0.7904
32x32 Random 33.3% of Negative Feedback Pairings 183,306 68.64% 0.541
@@ 4.5 Variations on Subnetwork Components
Name Trainable Params Count seqCIFAR10
@@
16x32 Global Workspace Included 22,538 57.92% 1.0921
24x32 Global Workspace Included 34,313 62.59% 0.8161
24x32 Global Workspace Repetition 34,313 61.52% 0.8644
24x32 Global Workspace Selection 34,313 62.32% 0.841
Size 64 Global Workspace With 24x32 Subnets 60,810 63.72% 0.6806
32x32 Global Workspace 46,090 62.5% 0.7904
Size 128 Global Workspace With 32x32 Subnets 147,210 65.57% 0.5532
Table 2: Performance of Global Workspace Sparse Combo Nets with different central module settings. The same stats as in Table 1 are presented here, now for an expanded set of global workspace experiments. This includes the initial tests where an otherwise normal node in the Sparse Combo Net was made to be the GW, alongside any repetitions run (Figure S2). Additionally, two new trials (highlighted rows) are reported – a size 64 central module connected to a $24\times 32$ set of contributing subnetworks initialized as before, and a size 128 central module connected to a $32\times 32$ set of contributing subnetworks initialized as before. In both cases the separate global workspace subnetwork was initialized using a $1\%$ sparsity level and weights between 5-5 and $5$. The larger GW subnetworks both resulted in test accuracy improvements, though they had notably larger size in terms of trainable parameters, and were perhaps slightly overfitted when comparing their training losses to those found in Table 1.
Note that while we used the notation $p\times n$ to refer to different size Sparse Combo Nets through most of this text for convenience, there is of course no requirement for network sizes to be the same, and subnetworks of a variety of sizes can be just as readily described within the existing model definition. Overall, the Size 128 Global Workspace with $32\times 32$ Subnets matched the performance of the best network from [1], and reached this accuracy within the first 200 epochs of training. Both networks had a similar total number of trainable parameters (Tables 2 and S1), though the GW network had many more units and thus has the greater potential for expansion to other applications. The size 128 GW network was also quite resistant to post-training subnetwork ablation (Figure S4), unlike the $16\times 32$ Sparse Combo Net with all-to-all negative feedback introduced previously [1] was (Figure 2).
@@ 5.1 Future Directions
Furthermore, we will investigate combination properties of networks that maintain functionality, rather than merely preserving stability. Such networks could include bespoke recurrent models, such as combinations of winner-take-all (WTA) networks, which are also contracting—i.e., their output dynamics depend on and stabilize with respect to their inputs [52]. Ensembles of recurrent neural networks pre-trained on individual tasks may also be considered. Predictive coding, a form of negative feedback, may be employed to learn feedback hierarchies—potentially using RNNs or similar architectures—and can also be analyzed with respect to modular stability [36, 53]. Additionally, we will consider the role of a reward signal in guiding network dynamics and maintaining functional integration.
@@
Note that task input at each timestep was first passed through a (trainable) linear input layer, represented by a weight matrix of size 3 (for RGB CIFAR pixels) by $N$, where N=pnN=p*n is the total number of units in the $p\times n$ Combo RNN. On the flip side, the outputs of the Combo RNN neurons at the final timestep in the sequence were used for prediction via passing them through a (trainable) linear output layer. The output layer was represented by a matrix of size $N$ by 10, with 10 being the number of different classes contained in the CIFAR10 dataset. This allowed for computation of cross entropy loss during training based on the true class of each input. At test time, the output index with the maximum value was selected as the predicted label.
@@
Name Stable RNN? Params sMNIST
@@
LSTM [51] 68K 97.3% 92.7% 59.7%
Transformer [54] 0.5M 98.9% 97.9% 62.2%
Antisymmetric [51] ? 36K 98% 95.8% 58.7%
Sparse Combo Net 130K 96.85% (4)
@@
As described in the main text, all experiments were run using the open sourced Sparse Combo Net code from [1] on Google Colab Pro+. Methodology for nonlinear subnetwork initialization and corresponding metric calculation was thus identical to the best performing trials of [1], as was the base model definition and general training loop. These settings were the same for all trials described in the main text, unless noted otherwise there. Across the whole paper (and all of [1]), Sparse Combo Net was run with $\tau=1$ and $dt=0.03$, such that after each timestep (i.e. after each pixel was presented), the current state was updated to be the previous state +0.03f(x)+0.03*f(x), where $f(x)=-x+\mathbf{W}*\phi(x)+\mathbf{L}*x+u(t)$ with $x$ corresponding to the previous state of the neurons and $u(t)$ corresponding to the values returned by the linear input layer when given the pixel that was just input to the larger network. $\mathbf{W}$ and $\mathbf{L}$ are of course the intra- and inter- subnetwork weights of (overall ”RNN of RNNs” dynamics) respectively, so they are static throughout this process of evaluating an individual image sequence. Recall that before each new sequence, network starting state was reset.
@@
Like discussed within Table 2, we did use different initialization settings for the global workspace subnetwork when experimenting with differing sizes for this central module. The main difficulty in comparing subnetworks of different sizes for the Sparse Combo Net in general is that sparsity levels cannot be fairly compared – for a larger nonlinear RNN subnetwork it may be impossible to meet the Theorem 1 condition of [1] without either increasing sparsity or greatly decreasing allowed entry magnitude, but for a smaller nonlinear RNN subnetwork a highly sparse structure might mean that the majority of units have 0 non-zero weights and thus cannot meaningfully contribute. Indeed, there was a point at which increasing subnetwork sparsity decreased the performance of Sparse Combo Nets [1]. As such, we used different sparsity level for initialization of those larger global workspace subnetworks in our GW size experiments. Specifically, we tested a GW with 64 units attached to 24 typically-defined 32 unit subnetworks, and a GW with 128 units attached to 32 typically-defined 32 unit subnetworks. In both cases, the GW subnetwork was initialized with a $1\%$ sparsity level and entry magnitude between 5-5 and $5$. The networks were otherwise trained as before, and their performance on seqCIFAR10 (Table 2) was then compared to similarly sized Global Workspace Sparse Combo Nets from the previous trials, which would have GW subnetwork of size 32 with $3.3\%$ sparsity level and entry magnitude between 6-6 and $6$ (as was used for the Sparse Combo Nets more broadly in [1]).
One additional technical consideration that may warrant further testing in different network topologies is the initialization scheme for inter-area connections. Like [1], we sample each entry from a normal distribution with mean of 0 and standard deviation $\frac{1}{\sqrt{n}}$ where $n$ is the number of neurons in each subnetwork of a $p\times n$ "RNN of RNNs". In the case where different subnetworks have different sizes, this instead uses the mean number of neurons per subnetwork. With a smaller number of between-subnetwork connections (as in the global workspace topology), one might consider allowing nonzero entries to be initialized to higher entry magnitudes. While in the present paper we have made the most direct comparison possible, this adjustment could allow advantages of the more sparsely connected negative feedback architectures to be even better utilized [^1] For a more extended set of relevant future experimental directions than provided in this paper, see Ennis [45]..
@@
Name Trainable Params Count seqCIFAR10
@@
37,258 61.07% 0.93
Size 64 Global Workspace With 24x32 Subnets 60,810 63.72% 0.6806
2x32 Global Workspace With 24x32 Subnets 61,834 63.31% 0.7529
Size 128 Global Workspace With 32x32 Subnets 147,210 65.57% 0.5532
4x32 Global Workspace With 32x32 Subnets 153,354 64.75% 0.6546